# DocReward: A Document Reward Model for Structuring and Stylizing Junpeng Liu1 Yuzhong Zhao211footnotemark: 1Bowen Cao1Jiayu Ding3Yilin Jia4Tengchao Lv5 Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5 Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei5 1CUHK 2UCAS

### v1

@@ Abstract
This gap arises mainly from the absence of suitable reward models to guide agentic workflows toward producing documents with stronger structural and stylistic quality.
@@
We construct a multi-domain dataset DocPair of 117K paired documents, covering 32 domains and 267 document types, each including a high- and low-professionalism document with identical content but different structure and style. This enables the model to evaluate professionalism comprehensively, and in a textual-quality-agnostic way.
@@
To assess the performance of reward models, we create a test dataset containing document bundles ranked by well-educated human evaluators.
Notably, DocReward outperforms GPT-4o and GPT-5 in accuracy by 30.6 and 19.4 percentage points, respectively, demonstrating its superiority over baselines.
In an extrinsic evaluation of document generation, DocReward achieves a significantly higher win rate of 60.8%, compared to GPT-5’s 37.7% win rate, demonstrating its utility in guiding generation agents toward producing human-preferred documents.
@@ 1 Introduction
Figure 1: DocReward automatically assesses document professionalism according to their structure and style, assisting existing agentic workflows for more professional document generation (left). It outperforms GPT-5 by 19.4% in human preference accuracy (right).
Recent advances in agentic workflows have automated many complex tasks, such as code generation (Peng et al., 2023; Cherny & Anthropic, 2025; Hong et al., 2024), image generation (comfyanonymous, 2025), visual understanding (Zheng et al., 2025; Marsili et al., 2025), math reasoning (Yan et al., 2025), and travel planning (Xie et al., 2024). A key focus of agentic workflows is the production of professional documents, including works like deep research (OpenAI, 2025a; Liang et al., 2025; Qwen, 2025) and technical documentation generation (Dvivedi et al., 2024). However, existing research about professional document generation primarily focuses on improving textual quality, neglecting the importance of visual structure and style, both of which play crucial roles in shaping document professionalism.
A well-organized structure helps readers navigate the material smoothly, while a consistent style makes the content more readable and engaging. Together, these aspects help convey information more clearly and effectively.
The neglect of structure and style mainly stems from the lack of suitable reward models, which are capable of guiding agentic workflows to produce documents with more professional structure and style.
To address this, we propose DocReward, a Document Reward Model, specialized in assessing document professionalism in structure and style, as shown in Figure 1.
@@
To achieve both comprehensiveness and textual-quality-agnosticism, we construct a multi-domain dataset, DocPair, consisting of 117K paired documents, covering 32 domains and 267 document types, with each pair consisting of a high-professionalism sample and its low-professionalism counterpart. The paired documents share identical content but differ in structure and style.
The construction of DocPair consists of three phases:
1) Curating High-Quality Professional Documents. We curate a set of high-quality documents with strong professionalism in structure and style, from various domains (, government, education, science, etc.)
2) Expanding Source Documents via Agents. Next, we extract both the textual content and the rendered pages of the source documents. Subsequently, multiple generation agents are prompted to produce a new document that preserves the textual content of the original and adheres to appropriate structure and style.
3) Ranking Documents. When comparing a source document with its generated counterparts, the original human-authored version is always preferred. In other cases, we use the original professional document as a reference and employ GPT-5 (OpenAI, 2025b) to rank document bundles by their structural and stylistic professionalism.
Based on the constructed dataset, we train DocReward to take rendered document pages as inputs and output a score reflecting the document’s professionalism in structure and style.
The predicted scores of paired documents are optimized using the Bradley-Terry loss (Bradley & Terry, 1952; Ouyang et al., 2022), which penalizes violations of the annotated order.
To demonstrate the superiority and utility of DocReward, we perform both intrinsic and extrinsic evaluations. For intrinsic evaluation, we create a test dataset of 473 human-annotated pairs across multiple document domains. Each pair is annotated by well-educated human evaluators, who assess the professionalism of the paired documents’ structure and style.
Notably, as shown in Figure 1 (right), DocReward outperforms GPT-4o (Hurst et al., 2024) and GPT-5 (OpenAI, 2025b) by 30.6 and 19.4 percentage points, respectively, in accuracy on the test dataset, demonstrating its superiority over existing approaches.
For extrinsic evaluation, in a human evaluation of document generation, DocReward as a reward model achieves a significantly higher win rate of 60.8%, compared to GPT-5’s 37.7%. This demonstrates its ability to guide generation agents in producing human-preferred documents, making it a valuable tool for improving document generation.
The contributions of this paper are summarized as follows:
@@
We propose DocReward, a Document Reward Model specialized in assessing document professionalism in terms of structure and style.
@@
To equip DocReward with comprehensiveness and textual-quality-agnosticism, we construct a multi-domain dataset DocPair, consisting of 117K paired documents across 32 domains and 267 document types. This enables the model to evaluate professionalism in structure and style comprehensively and independently of inherent textual content quality.
We propose a human-annotated test dataset for assessing document professionalism in structure and style. Experimental results show that DocReward outperforms strong baselines including GPT-5. Furthermore, human evaluation of document generation demonstrates its effectiveness in improving document generation quality.
2 Task Formulation
@@
A document’s professionalism is determined by its textual content, structure, and style. Although large language models excel at evaluating textual quality, they are limited in assessing structure and style. To bridge this gap, we develop reward models tailored to these dimensions to advance agentic workflows in producing documents with more professional structure and style.
In this section, we formulate the task and provide a clear definition of its objectives.
@@
Let denote a set of documents, where each document consists of textual content and rendered images . The document reward model assigns scores to documents that share the same textual content, such that the scores reflect their structural and stylistic quality. This process is formalized as follows:
@@
Structure: proper use of white space, appropriate margins, clear section breaks, well-structured text alignment, adequate paragraph spacing, proper indentation, inclusion of page headers and footers, and logical, coherent organization of content.
Style: appropriate font choices (type, size, color, readability), clear heading styles, effective use of emphasis (bold, italics), bullet points, numbering, and consistent formatting.
By optimizing based on these factors, we obtain a reward model capable of assessing the structural and stylistic professionalism in a comprehensive and textual-quality-agnostic way.
@@ 3 DocReward
We propose DocReward, a reward model specializing in assessing the structural and stylistic professionalism of documents. DocReward is trained on DocPair, a diverse dataset of 117K document pairs (Section 3.1), and is optimized with a preference-based objective for structural and stylistic assessment (Section 3.2). The following sections detail the data construction pipeline and model design.
@@ 3.1 Data Construction
The overall process results in DocPair, a dataset comprising 117K document pairs, covering 32 domains and 267 document types. The construction procedure is detailed step by step below:
Curating High-Quality Professional Documents.
As illustrated in Figure 2 (top), we first curate a corpus of human-authored Microsoft Word documents that spans both highly formal institutional writing and everyday professional communication. We draw on two complementary sources:
Government and institutional corpora: GovDocs1 (Garfinkel et al., 2009) and NapierOne (Davies et al., 2022). GovDocs1 is a publicly available collection compiled from U.S. government (.gov) websites, including policy reports, administrative forms, statistical reports, public guidance, and meeting minutes, etc. NapierOne is a modern, comprehensive document dataset sourced from a wide range of public institutional materials and common office documents. These corpora provide authoritative, consistently professional exemplars of document structure and style.
Web document corpus: We also draw from a diverse set of documents discovered in the CommonCrawl repository111https://commoncrawl.org/. This corpus captures a broad range of real-world professional documents from business, education, nonprofit, healthcare, and other sectors, such as proposals, syllabi, newsletters, technical manuals, and policy briefs. It substantially enhances structural and stylistic diversity across professional genres.
To ensure suitability for reward-model training, we apply a light-weight preprocessing and filtering pipeline before data construction. First, all files are converted to DOCX format to enable programmatic access and modification via python-docx222https://python-docx.readthedocs.io/en/latest/. Next, we discard extreme or malformed cases (exceeding 20 pages, files larger than 1 MB dominated by images, and files smaller than 10 KB with trivial content). To efficiently reduce residual noise, we employ GPT-5 as a rigorous automated heuristic to flag poor structure/style on a scale; documents scoring above 8 are retained. A manual inspection of 200 randomly sampled retained documents confirms that this automated filter preserves high-quality professional samples.
Finally, we analyze the distribution of domains and document types to assess coverage. The filtered collection spans 32 domains (e.g., government, education, nonprofit, medical, scientific, legal, business, academic, technical) and over 267 document types (e.g., job descriptions, government forms, policy documents, meeting minutes, press releases, course syllabi). The top 10 domains and top 30 document types are shown in Figure 4 and Figure 4, respectively, demonstrating both breadth and diversity. These high-quality, professional documents form the foundation for constructing subsequent document bundles and comparison pairs.
@@
As shown in Figure 2 (middle), to obtain documents with the same textual content but different structure and style, we construct two types of agents to synthesize new documents given the textual content (and rendered pages) of the source documents. To further increase the diversity of the synthesized documents, each agent can be empowered by different LLMs. The two proposed agents are detailed as follows:
Textual Content to Document.
The textual content is first extracted from the source documents, discarding all formatting, styling, and layout information. Then, advanced generation agents (, GPT-4o, OpenAI o1 (OpenAI, 2024), Claude Sonnet 4 (Anthropic, 2025), and GPT-5) are used to synthesize DOCX documents via python-docx. This process simulates the real-world task of generating professionally structured and styled documents from plain text.
Refinement for Better Structure and Style.
To further improve the structure and style of synthesized documents, we refine them by comparing with the original human-authored documents in terms of structure and style. The refinement process consists of two stages:
1) Generation agents are provided with the python-docx code, rendered pages, and structured textual representation of the synthesized document, along with the rendered pages of the original human-authored document, to generate a refinement plan.
2) Using this refinement plan, the agents modify the python-docx code to produce refined documents with better structure and style.
Since generation agents may omit textual content from the original documents, we remove any synthesized documents whose textual content deviates significantly from that of the original human-authored one. The remaining synthesized documents are then grouped with their originals to facilitate subsequent processing. For the details and prompts of this phase, please refer to Appendix A.3 and Appendix A.4.
@@
As shown in Figure 2 (bottom), the documents within the same group share identical textual content and are organized into pairs. The annotation task is to assess the relative professionalism in terms of structure and style for each pair, which is carried out under the following two cases:
Real v.s. Synth. If any sample in the pair is from the human-authored professional documents 3.1, it is directly designated as the preferred (winner).
Synth v.s. Synth. When both samples in the pair are generated by agents, we prompt GPT-5 with a document triplet , where the human-authored professional document is used as a reference to decide which synthetic sample is preferred. GPT-5 achieves an average accuracy of 92.5% on a human-annotated evaluation set consisting of 120 pairs in our preliminary test, demonstrating that the triple-wise annotation method is reliable and well-aligned with human judgment.
The prompt is presented in Appendix A.4.
The two types of annotations are both guided by human-authored professional documents, and serve complementary purposes: “Real vs. Synth” pairs steer agentic workflows toward human-level document generation, while “Synth vs. Synth” pairs promote self-refinement.
The data statistics of the constructed dataset, , DocPair, are shown in Table 1.
@@
Doc. Pairs
Total
Real Synth
Synth Synth
@@ 3.2 Model Structure and Optimization
We adopt Qwen-2.5-VL (Bai et al., 2025) as the base model due to its advanced native multi-image input capabilities, which allow for a more comprehensive analysis of multi-page documents. An N-page document is converted into N images, which are then input into the model. A regression head is added to predict a scalar score on top of the output hidden states. More implementation details are presented in Appendix A.2.
We optimize DocReward using the Bradley-Terry (BT) loss, which is specifically designed for learning from pairwise preferences.
Specifically, let and be the rendered pages of the preferred (winner) and those of the less preferred (loser) in a paired comparison, respectively, then, the DocReward (formatted as ), takes in the rendered pages of each document and outputs scores, separately, which are supervised with the following objective:
where is the sigmoid function, defined as . This objective encourages the model to assign a higher score to the preferred document compared to the less preferred one.
@@ 4 Experiments
We conduct experiments to evaluate the effectiveness of DocReward in assessing both structural and stylistic professionalism of documents. This section includes evaluation dataset annotation, quantitative comparisons with strong baselines, extrinsic evaluation of document generation, and qualitative analyses.
4.1 Evaluation Dataset Collection and Human Annotation
@@
A subset of the curated documents in Section 3.1 is set aside as evaluation documents. To diversify the evaluation dataset, we consider the following six types of documents using the method described in Section 4. Four of them are obtained via the Textual Content to Document agent, which generates DOCX documents using different LLMs (, GPT-4o, OpenAI o1, Claude Sonnet 4, and GPT-5). One type comes from the Refinement for Better Structure and Style agent, where GPT-5 is employed to refine synthesized documents. The last type consists of the original human-authored documents. Together, these six types constitute the origins of samples in our evaluation dataset.
@@
For each group of documents sharing the same content but differing in structure and style, human experts meticulously rank their quality based on structure and style. To facilitate model evaluation, these ranked relationships are converted into a total of 473 comparison pairs, each consisting of two documents and a binary label indicating the preferred one. To ensure the quality of human annotation, two highly educated annotators annotate the same subset of documents; then, we calculate the overall percentage of decisions on which the annotators agreed, which is 91.6%. This demonstrates the good agreement between annotators.
4.2 Baselines and Evaluation Settings
We evaluate our approach against several strong language models, including GPT-4o, Claude Sonnet 4, and GPT-5. Two evaluation settings are considered: pairwise and pointwise. In the pairwise setting, the model receives the rendered pages of two documents and is instructed to predict which document exhibits superior structure and style. In the pointwise setting, the model is provided with the rendered pages of a single document and assign a scalar score for structure and style without any reference document. The evaluation metric is accuracy, defined as the proportion of predictions that correctly match human annotations in the evaluation dataset.
Table 2: Accuracy of Models on the proposed evaluation dataset. ’Real vs. Synth’ represents evaluation pairs where a human-authored document is compared against a document generated by an agent. ’Synth vs. Synth’ represents evaluation pairs where two agent-generated documents are compared.
Superior Performance of DocReward over Baselines.
As presented in Table 2, on the human-annotated evaluation dataset, DocReward-3B and DocReward-7B, achieve substantial improvements over strong baselines including GPT-4o, Claude Sonnet 4, and GPT-5.
In particular, DocReward-7B achieves an overall human preference accuracy of 89.22% , 19.45 points higher than the strongest closed-source baseline (GPT-5, 69.77%).
In the critical “Real vs. Synth” setting, DocReward-7B achieves 97.42%, indicating near-perfect alignment with human judgments when distinguishing professional human-authored documents from synthetic ones. Even in the more challenging “Synth vs. Synth” setting, DocReward-7B maintains 78.22%, substantially higher than GPT-5 (64.85%).
These results demonstrate that DocReward effectively captures structural and stylistic quality signals that existing LLMs overlook.
Position Bias in Pairwise Baselines.
To assess potential order effects in pairwise evaluation, we tallied how often each baseline model selected the first versus the second document as the preferred option. As shown in Table 4. The analysis reveals that GPT-4o and Claude Sonnet 4 exhibit a noticeable position bias, with a consistent tendency to favor the second document in the pair. In contrast, GPT-5 shows almost no such bias, producing balanced selections across positions. This finding highlights an important caveat for developing future pairwise reward models: even when document order is randomized, systematic position biases can distort evaluation results. Our proposed DocReward, being a pointwise model, does not suffer from such order effects, ensuring more stable and unbiased preference predictions.
Reward Models
Position 1
Position 2
GPT-4o
202
271
Claude Sonnet 4
189
284
GPT-5
240
233
Table 3: Position Preferred Times of Pairwise Baselines.
Reward Models

### v2

@@ Abstract
This gap stems mainly from a lack of effective reward models capable of guiding agents toward producing documents with high structural and stylistic professionalism.
@@
The model is trained under a textual-quality-agnostic framework to assess professionalism without being influenced by textual quality. To achieve this,
we construct a multi-domain dataset DocPair of 117K paired documents, covering 32 domains and 267 document types, each comprising a high- and low-professionalism document with identical content but different structure and style.
This setup enables the model to evaluate professionalism comprehensively and independently of textual quality.
@@
On a manually annotated benchmark, DocReward outperforms GPT-5 by 14.6 percentage points in accuracy.
Extrinsic RL experiments further validate its effectiveness in guiding professional document generation.
DocReward: A Document Reward Model for Structuring and Stylizing
Junpeng Liu1††thanks: Equal contribution. Work done during internship at Microsoft Research.
Yuzhong Zhao211footnotemark: 1
Bowen Cao1
Jiayu Ding3
Yilin Jia4
Tengchao Lv5Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei51CUHK
2UCAS
3XJTU
4UMich
5Microsofthttps://aka.ms/GeneralAI
@@ 1 Introduction
Figure 1: DocReward automatically assesses document professionalism according to their structure and style, assisting existing agentic workflows for more professional document generation (left). It outperforms GPT-5 by 14.6% in human preference accuracy (right).
Recent advances in agentic workflows have automated various complex tasks, such as code generation (Peng et al., 2023; Cherny and Anthropic, 2025; Hong et al., 2024), image generation (comfyanonymous, 2025), visual understanding (Zheng et al., 2025; Marsili et al., 2025), math reasoning (Yan et al., 2025), and travel planning (Xie et al., 2024). A key focus of agentic workflows is the generation of professional documents, including works like deep research (OpenAI, 2025b; Liang et al., 2025; Qwen, 2025) and technical documentation generation (Dvivedi et al., 2024). However, existing research on professional document generation primarily focuses on improving textual quality, often neglecting the visual structure and style—both of which are crucial for document professionalism.
A well-organized structure facilitates seamless navigation for readers, while a consistent style makes the content more readable and engaging. Together, these aspects help convey information more clearly and effectively.
The neglect of structure and style mainly stems from the lack of effective reward models, which are capable of guiding agentic workflows toward producing documents with professional structure and style.
@@
To achieve this, we propose DocReward, a Document Reward Model specialized in assessing document professionalism in structure and style, as illustrated in Figure 1.
The model is trained under a textual-quality-agnostic framework. Specifically,
we construct DocPair, a multi-domain dataset of 117K document pairs across 32 domains and 267 types. Each pair comprises a high-professionalism sample and its low-professionalism counterpart; these documents share identical content but differ in structure and style.
The construction of DocPair involves three phases:
1) Curating High-Quality Professional Documents. We curate a collection of high-quality documents characterized by professional structure and style from diverse domains, such as government, education, and science.
2) Expanding Source Documents via Agents. Next, we extract the textual content and employ multiple generation agents to re-produce documents that preserve the original textual content while exploring diverse structural and stylistic variations.
3) Ranking Documents. The ranking labels are determined by a combination of human-verified heuristics and an oracle-based annotation method. This provides a scalable way to capture elements of structural and stylistic professionalism.
Based on the constructed dataset, we train DocReward to take rendered document pages as input and output a score reflecting the document’s professionalism in structure and style.
The predicted scores of paired documents are optimized using the Bradley-Terry loss (Bradley and Terry, 1952; Ouyang et al., 2022), which penalizes violations of the annotated ranking order.
To demonstrate the superiority and utility of DocReward, we conduct both intrinsic and extrinsic evaluations.
For the intrinsic evaluation, we establish a benchmark DocPairBench of 1443 human-annotated pairs across multiple domains. Human annotators ranked each pair based on the professionalism of the documents’ structure and style. Notably, as shown in Figure 1 (right), DocReward outperforms GPT-4o (Hurst et al., 2024) and GPT-5 (OpenAI, 2025a) by 27.4 and 14.6 percentage points, respectively, in accuracy on DocPairBench, demonstrating its superiority over existing baselines.
For the extrinsic evaluation, we evaluate DocReward through two complementary experiments.
1) Best-of-N. DocReward is used as a re-ranking model for improving agentic workflow without changing the agent itself. Human evaluation reveals that DocReward as a reward model achieves a significantly higher win rate of 60.8%, compared to GPT-5’s 37.7%.
2) Reinforcement Learning. We further demonstrate the utility of DocReward as the reward model for reinforcement-learning of both open- and closed-source models. This integration improves the document generation performance of Qwen2.5-Coder (Hui et al., 2024) and GPT-4o regarding structure and style.
Our main contributions are as follows:
We propose a textual-quality-agnostic framework for document reward modeling. By decoupling structure and style from textual content, it enables the reward model DocReward to accurately capture complex structural and stylistic elements.
@@
DocPair: A large-scale multi-domain dataset comprising 117K document pairs across 32 domains and 267 document types, designed to equip DocReward with comprehensiveness and textual-quality-agnosticism.
Comprehensive experiments demonstrate that DocReward not only outperforms GPT-5 in assessing structure and style, but also serves as an effective reward model in reinforcement learning for agentic workflows.
2 Textual-Quality-Agnostic Framework
@@
Document professionalism is characterized by its textual content, structure, and style. Although large language models excel at evaluating textual quality, they are limited in assessing structure and style.
To bridge this gap, we propose a textual-quality-agnostic framework that trains reward models to evaluate professionalism independently of textual quality.
In this section, we formulate the framework and define its objectives.
@@
Let denote a set of documents, where each document consists of textual content and rendered images . The document reward model assigns scores to documents that share the same textual content, such that these scores reflect their structural and stylistic quality. This process is formalized as follows:
@@
Essentially, the term “” ensures that the model evaluates professionalism in a textual-quality-agnostic manner, as it processes identical textual content from paired documents.
@@
1) Structure: Proper use of white space, appropriate margins, clear section breaks, well-structured text alignment, adequate paragraph spacing, proper indentation, inclusion of page headers and footers, and logical, coherent organization of content.
2) Style: Appropriate font choices (type, size, color, readability), clear heading styles, effective use of emphasis (bold, italics), bullet points, numbering, and consistent formatting.
By optimizing based on these factors, we obtain a reward model capable of assessing structural and stylistic professionalism in a comprehensive and textual-quality-agnostic way.
@@ 3 DocReward
Based on the proposed textual-quality-agnostic framework, we train DocReward, a reward model specializing in assessing the structural and stylistic professionalism of documents. DocReward is trained on DocPair, a diverse dataset of 117K document pairs (Section 3.1), and is optimized with a preference-based objective for structural and stylistic assessment (Section 3.2). The following sections detail the data construction pipeline and model design.
@@ 3.1 Data Construction
The construction procedure is detailed step by step below:
Curating High-Quality Professional Documents. As illustrated in Figure 2 (top), we first curate a corpus of human-authored Microsoft Word documents that spans both highly formal institutional writing and everyday professional communication. We draw on two complementary sources:
1) Government corpora: GovDocs1 (Garfinkel et al., 2009) and NapierOne (Davies et al., 2022) are authoritative document collections sourced from government and public institutions, covering reports, forms, guidelines, and other professional materials with consistent structure and style.
2) Web document corpus: CommonCrawl111https://commoncrawl.org/, covering business, education, healthcare, etc, enriching structural and stylistic diversity.
To ensure high quality, we apply a preprocessing and filtering pipeline before data construction. First, all files are converted to DOCX format to enable programmatic access and modification via python-docx222https://python-docx.readthedocs.io/. Next, we discard extreme or malformed cases (exceeding 20 pages, files larger than 1 MB dominated by images, and files smaller than 10 KB with trivial content). To efficiently reduce residual noise, we employ GPT-5 as a rigorous automated heuristic to flag poor structure/style on a scale; documents scoring above 8 are retained. A manual inspection of 200 randomly sampled retained documents confirms that this automated filter preserves high-quality professional samples.
The domain distributions of the filtered corpus are shown in Figure 3. The corpus spans 32 domains and over 267 document types, demonstrating substantial breadth and diversity, and it provides a high-quality foundation for constructing subsequent document pairs.
@@
As shown in Figure 2(middle), to obtain documents with the same textual content but different structure and style, we construct two types of agents to synthesize documents given the textual content (and rendered pages) of the source documents. To further increase the diversity of the synthesized documents, each agent can be empowered by different LLMs. The two agents are detailed as follows:
1) Textual Content to Document.
The textual content is first extracted from the source documents, discarding all formatting, styling, and layout information. Then, advanced generation models (, GPT-4o, OpenAI o1 (OpenAI, 2024), Claude Sonnet 4 (Anthropic, 2025), and GPT-5) are used to synthesize DOCX documents via python-docx.
2) Refinement for Better Structure and Style.
To further improve the structure and style of synthesized documents, we refine them by comparing them with the original human-authored documents in terms of structure and style. The refinement process consists of two stages:
1) Agents are provided with the python-docx code, rendered pages, and structured textual representation of the synthesized document, along with the rendered pages of the original human-authored document, to generate a refinement plan.
2) Based on the plan, the agents modify the python-docx code to produce refined documents with improved structure and style.
The synthesized documents are grouped with their originals to facilitate subsequent processing. Please refer to Appendix A.3 and A.6 for details.
@@
As shown in Figure 2 (bottom), the collected documents within each group share identical textual content and are organized into pairs. The annotation task is to assess the relative professionalism in terms of structure and style for each pair, which is carried out under the following two cases:
1) Real vs. Synth.
When comparing an original document with its agent-generated counterparts, the human-authored version is designated as the winner. This heuristic is grounded in our preliminary human inspection, which indicates that current state-of-the-art models (e.g., GPT-5, Claude Sonnet 4) produce documents with structure and stylistic quality inferior to high-quality human-authored documents curated by the rigorous filtering pipeline (described in 3.1).
2) Synth. vs. Synth. For pairs where both documents are agent-generated, manual annotation is not scalable. We therefore employ a closed-source model as a proxy for human judgment. To mitigate its intrinsic biases, we adopt an oracle setting: the model is provided with a document triplet , using the human-authored as a reference anchor. This transforms the annotation from a subjective preference task into an objective similarity matching against the ground-truth document (see subsection A.6).
We validated this approach on 120 pairs, where it demonstrated 92.5% alignment with human experts, confirming its effectiveness.
Note that the above strategy is cost-effective and necessary for large-scale training data construction, while the test split, DocPairBench, is constructed through strict human annotation. The data statistics of DocPair are shown in Table 1.
@@
Pairs
@@ 3.2 Model Structure and Optimization
We use Qwen-2.5-VL (Bai et al., 2025) as the base model, which takes a multi-page document as input and outputs a scalar score via an added regression head (details provided in Appendix A.2).
Given a preference pair consisting of a winner and a loser , the reward model assigns scores to both documents and is optimized using the Bradley–Terry (BT) loss:
where . This objective encourages higher scores for preferred documents.
@@ 4 Experiments
We first present DocPairBench (Section 4.1) and then conduct intrinsic and extrinsic evaluations as shown in Figure 4 (Section 4.2 and 4.3).
Figure 4: Three evaluation settings for DocReward.
4.1 DocPairBench
@@
A subset of the curated documents in Section 3.1 is set aside as evaluation documents. To diversify the evaluation samples, we consider the following six types of documents using the method described in Section 3. Four of them are obtained via the Textual Content to Document agent, which generates DOCX documents using different LLMs (, GPT-4o, OpenAI o1, Claude Sonnet 4, and GPT-5). One type comes from the Refinement for Better Structure and Style agent, where GPT-5 is employed to refine synthesized documents. The last type consists of the curated human-authored documents. Together, these six types constitute the origins of samples in our benchmark.
@@
For each set of documents sharing the same content but differing in structure and style, human experts meticulously rank their quality based on structure and style. To facilitate model evaluation, these ranked relationships are converted into a total of 1443 comparison pairs, each consisting of two documents and a binary label indicating the preferred one. To ensure the quality of human annotation, we evaluate annotation consistency among human annotators using Cohen’s Kappa and observe a high agreement of 83.4.
The details of annotation protocol and reliability validation are presented in Appendix A.4.
Table 2: Accuracy of Models on the DocPairBench benchmark across ten domains: Government (Gov.), Education (Edu.), Non-Profit Organization (NPO), Medical (Med.), Science (Sci.), Legal (Leg.), Business (Bus.), Academic (Acad.), Technology (Tech.), and Other (Oth.).
As presented in Table 2, we evaluate all methods under two settings, namely Pairwise and Pointwise. In the pairwise setting, models are given a document pair and asked to select the one with better structure and style, while in the pointwise setting, documents are scored independently. Performance is measured by accuracy with respect to human annotations. On DocPairBench, DocReward-3B and DocReward-7B achieve substantial improvements over strong baselines including Qwen2.5-VL, GPT-4o, Claude Sonnet 4, and GPT-5. In particular, DocReward-7B attains an overall human preference accuracy of 82.3 percent, surpassing the strongest closed source model (GPT-5) by 14.6 points. These results indicate that DocReward captures structural and stylistic quality signals that existing language models often overlook.
4.3 Improving Document Generation with DocReward
To demonstrate the utility of DocReward, we conduct two complementary experiments.
Rewards